# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ Appendix B PROOF OF MAIN TEXT RESULTS
$KL_{Y}\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)$ $=KL\left(\widehat{P}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P^{(\omega(\mathbf{i}))}({Y},{\mathbf{Z}})\right)-\mathbb{E}_{y\sim\widehat{P}^{(\mathbf{i})}\left(Y\right)}\left[KL_{Y}\left(\widehat{P}^{(\mathbf{i})}\left({\mathbf{Z}}|{Y}=y\right)||{P}^{(\omega(\mathbf{i}))}\left({\mathbf{Z}}|{Y}=y\right)\right)\right]$
$\leq KL\left(\widehat{P}^{(\mathbf{i})}(\widehat{Y},{\mathbf{Z}})||P^{(\omega(\mathbf{i}))}({Y},{\mathbf{Z}})\right)=\mathcal{L}_{cons}\,.$
@@
B.2 Proof of Proposition 3.3
See 3.3
In addtion to the proof of this proposition, we are going to show that Assum. 3.2 can be replace by the following Assum. B.2 in this proposition, to yield the same result.
Assumption B.2.
$P(\mathbf{i}_{\pi(1)})$ is a mixture of $\pi(1)+1$ distinct diracs, including one at zero. The remaining compenents have constant value zero.
@@
We exploit the positive definiteness of the KL loss and its continuity with respect to $\mathbf{i}$. Since the variables are jointly Gaussian, continuity is obvious from the analytical expression of the KL for Gaussian variables and continuity of the shift operation applied to the parameters of the Gaussian. We exploit the cause-mechanism decomposition and the lower-bound by $\mathcal{L}_{cons}$ to progressively identify necessary conditions on parameters to have $\mathcal{L}_{cons}=0$ and finally check those conditions are sufficient.
Let $N_{0}$ be the size of $\pi(0)$ and $N_{1}$ be the size of $\pi(1)$. Because of the SCM assumption, the causal graph is a DAG, such that variables can be ordered without loss of generality such that $A$ (and thus $A_{00}$ and $A_{11}$) is strictly upper triangular. This entails that $(I_{N}-A)$ is invertible (as an upper triangular matrix with non-zero diagonal coefficients). The low-level variables then satisfy
@@
where
$(I_{N}-A)^{-1}=\begin{bmatrix}(I_{N_{0}}-A_{00})^{-1}\,,&(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}\\ 0&(I_{N_{1}}-A_{11})^{-1}\end{bmatrix}$
leading to the pushforward high-level cause variable
$\widehat{Z}_{1}^{(\mathbf{i})}=\bm{\tau}_{1}^{\top}(I_{N}-A)^{-1}(\mathbf{U}+\mathbf{i})=\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}(\mathbf{U}_{\pi(1)}+\mathbf{i}_{\pi(1)})\,.$
Looking for the solutions satisfying $\mathcal{L}_{cons}=0$ entails that the must satisfy that for almost all $\mathbf{i}$, this push-forward distribution matches the learned high-level interventional distribution which satisfies
@@
According to Assum. 3.2, and since we assume $\Omega=\pi(1)$, the prior $P(\mathbf{i}_{\pi(1)})$ has density with respect to the Lebesgue measure with support including a neighborhood of $\mathbf{i}=\bm{0}$, by continuity of the KL divergence, a solution making the consistency loss vanish needs to have the KL divergence term vanish for $\mathbf{i}=\bm{0}$ (otherwise we could find a neighborhood of $\mathbf{i}=\bm{0}$ such that the KL does not vanish, by continuity of the KL divergence).
This vanishing of the KL divergence entails that its terms: the two unintervened densities, are equal, such that we get
$P^{(0)}({Z}_{1})=(\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1})_{\#}P(\mathbf{U})=\mathcal{N}(\bar{\bm{\tau}}_{1}^{\top}(I_{N}-A_{11})^{-1}\mu_{\mathbf{U}_{1}},\bar{\bm{\tau}}_{1}^{\top}(I_{N}-A_{11})^{-1}\Sigma_{\mathbf{U}_{\pi(1)}}(I_{N}-A_{11})^{-\top}\bar{\bm{\tau}}_{1})\,,$
which entails
$\sigma_{Z,1}^{2}=\bar{\bm{\tau}}_{1}^{\top}(I_{N}-A_{11})^{-1}\Sigma_{\mathbf{U}_{\pi(1)}}(I_{N}-A_{11})^{-\top}\bar{\bm{\tau}}_{1}$ (12)
@@
$\mu_{Z,1}=\bar{\bm{\tau}}_{1}^{\top}(I_{N}-A_{11})^{-1}\mu_{\mathbf{U}_{\pi(1)}}\,.$ (13)
For the same reasons, we can further match the interventional distributions in an open set included in the interior of the support of $P(\mathbf{i})$, such that for all $\mathbf{i}$ in this open set
$\mathcal{N}(\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}(\mu_{\mathbf{U}_{1}}+\mathbf{i}),\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\mathbf{U}_{1}}(I_{N_{1}}-A_{11})^{-\top}\bar{\bm{\tau}}_{1})\\ =\mathcal{N}(\bar{\tau}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\mu_{\mathbf{U}_{1}}+{\bm{\omega}}_{1}^{\top}\mathbf{i},\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\mathbf{U}_{1}}(I_{N_{1}}-A_{11})^{-\top}\bar{\bm{\tau}}_{1})\,.$
Indeed, otherwise the KL would not vanish in a neighborhood of non-zero measure and would contradict the fact that $\mathcal{L}_{cons}$ vanishes.
@@
$\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}(\mu_{\mathbf{U}_{\pi(1)}}+\mathbf{i}_{\pi(1)})=\bar{\tau}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\mu_{\mathbf{U}_{\pi(1)}}+{\bm{\omega}}_{1}^{\top}\mathbf{i}_{\pi(1)}\,,$
@@
$\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\mathbf{i}_{\pi(1)}={\bm{\omega}}_{1}^{\top}\mathbf{i}_{\pi(1)}\,.$
Since this equality between two affine functions of $\mathbf{i}_{\pi(1)}$ is valid on an open set of the vector space of $\mathbf{i}_{\pi(1)}$, these affine functions must be equal (we can reparameterize $\mathbf{i}$ to show that constants must match at the new origin and the linear maps must match on a basis of the space, so they are equal). This is valid if and only if, in addition to (12), (13),
${\bm{\omega}}_{1}=(I_{N_{1}}-A_{11})^{-\top}\bar{\bm{\tau}}_{1}\,,$ (14)
@@
Alternatively, we obtain the same result by replacing Assum. 3.2 by Assum. B.2. Indeed, as long as the collection of finite interventions vectors forms a rank $\pi(1)$ family, we can choose a subset of $\pi(1)$ such vectors $\{\mathbf{i}^{1}_{\pi(1)},\dots,\mathbf{i}_{\pi(1)}^{\pi(1)}\}$ such that it forms a linearly independent family. It can be used to build the matrix equality
$\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\left[\mathbf{i}^{1}_{\pi(1)},\dots,\mathbf{i}_{\pi(1)}^{\pi(1)}\right]={\bm{\omega}}_{1}^{\top}\left[\mathbf{i}^{1}_{\pi(1)},\dots,\mathbf{i}_{\pi(1)}^{\pi(1)}\right]$ (15)
where the matrix $\left[\mathbf{i}^{1}_{\pi(1)},\dots,\mathbf{i}_{\pi(1)}^{\pi(1)}\right]$ is invertible. By right-multiplying Eq. (15) by this inverse, we obtain Eq. (14) again.
We can move on to check the implication of consistency of the effect’s conditional. It entails for almost all of $\mathbf{i}$
@@
The left-hand side is obtained by using
@@
$Y=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\mathbf{U}_{\pi(0)}+\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}(\mathbf{U}_{\pi(1)}+\mathbf{i})$
@@
As we look for linear Gaussian high-level models, we assume $f$ is affine and parametrize it as $f(Z)=\alpha Z+\beta$. Then, under Assum. 3.2, equality of marginal distributions entails the following equality for all $\mathbf{i}_{\pi(0)}$ in an open neighborhood of 00 (otherwise $\mathcal{L}_{rel}\leq\mathcal{L}_{cons}$ would not vanish)
$\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\bm{\mu}_{\mathbf{U}_{0}}+\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}(\bm{\mu}_{\mathbf{U}_{1}}+\mathbf{i}_{\pi(1)})=\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}(\mu_{\mathbf{U}_{1}}+\mathbf{i}_{\pi(1)})+\beta+\bm{\mu}_{R_{0}}$
@@
$+\beta+\bm{\mu}_{R_{0}}=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\bm{\mu}_{\mathbf{U}_{\pi(0)}}\,.$
We can fix $\bm{\mu}_{R_{0}}$ to zero to avoid redundancy, such that
$\mu_{Y|Z=0}=\beta=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\bm{\mu}_{\mathbf{U}_{\pi(0)}}$ (16)
and consistency of non-zero shift interventions additionally entail for all $\mathbf{i}$ in the support of $P(\mathbf{i})$
$\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}\mathbf{i}_{\pi(1)}=\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\mathbf{i}_{\pi(1)}\,.$
This yields (since on can always choose a linearly independent family of vectors $\mathbf{i}_{\pi(1)}$ within the open neighborhood of zero for which this equality holds)
$\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}=\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\,.$
Then, right-multiplying by $(I_{N_{1}}-A_{11})$, we get
$A_{01}^{\top}(I_{N_{0}}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}=\alpha\bar{\bm{\tau}}_{1}\,.$ (17)
Similarly as above, the same conclusion can be drawn if we replace Assum. 3.2 by Assum. B.2.
We can finally check that the conditional distributions are matching. Let us first compute the covariance matrix of the low-level variables.
$cov(\mathbf{X})=\begin{bmatrix}(I_{N_{0}}-A_{00})^{-1}&(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}\\ \bm{0}&(I_{N_{1}}-A_{11})^{-1}\end{bmatrix}\Sigma_{U}\begin{bmatrix}(I_{N_{0}}-A_{00})^{-\top}&\bm{0}\\ (I_{N_{1}}-A_{11})^{-\top}A_{01}^{\top}(I_{N_{0}}-A_{00})^{-\top}&(I_{N_{1}}-A_{11})^{-\top}\end{bmatrix}\,.$
Because the exogenous covariance is diagonal, we get
$=\begin{bmatrix}\begin{matrix}(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}A_{01}^{\top}(I_{N_{0}}-A_{00})^{-\top}\\ +(I_{N_{0}}-A_{00})^{-1}\Sigma_{\pi(0)}(I_{N_{0}}-A_{00})^{-\top}\end{matrix}&(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}\\ \hline\cr(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}A_{01}^{\top}(I_{N_{0}}-A_{00})^{-\top}&(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}\end{bmatrix}\,.$
Then,

### v2

@@ Appendix C PROOF OF MAIN TEXT RESULTS
$\mathrm{KL}_{Y}\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)$ $=\mathrm{KL}\left(\widehat{P}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P^{(\omega(\mathbf{i}))}({Y},{\mathbf{Z}})\right)-\mathbb{E}_{y\sim\widehat{P}^{(\mathbf{i})}\left(Y\right)}\left[\mathrm{KL}_{Y}\left(\widehat{P}^{(\mathbf{i})}\left({\mathbf{Z}}|{Y}=y\right)||{P}^{(\omega(\mathbf{i}))}\left({\mathbf{Z}}|{Y}=y\right)\right)\right]$
$\leq\mathrm{KL}\left(\widehat{P}^{(\mathbf{i})}(\widehat{Y},{\mathbf{Z}})||P^{(\omega(\mathbf{i}))}({Y},{\mathbf{Z}})\right)=\mathcal{L}_{\mathrm{cons}}\,.$
@@
C.2 Proof of Proposition 3.4
See 3.4
@@
Part 1: 1-cause TCR.
Overview. We exploit the positive definiteness of the KL loss and its continuity with respect to $\mathbf{i}$. Since the variables are jointly Gaussian, continuity is obvious from the analytical expression of the KL for Gaussian variables and continuity of the shift operation applied to the parameters of the Gaussian. We exploit the cause-mechanism decomposition and the lower-bound by $\mathcal{L}_{\mathrm{cons}}$ to progressively identify necessary conditions on parameters to have $\mathcal{L}_{\mathrm{cons}}=0$ and finally check those conditions are sufficient.
Preliminaries. Let $N_{0}$ denote the size of $\pi(0)$ and $N_{1}$ be the size of $\pi(1)$. The SCM is assumed uniquely solvable (Definition 2.1), such that $\mathbf{x}=A\mathbf{x}+\mathbf{u}$ has a unique solution for almost all values of $\mathbf{u}$. Since $\mathbf{U}$ has full support, this implies that $I_{N}-A$ is invertible. The low-level variables then satisfy
@@
where, due to the block triangular form of $A$,
$(I_{N}-A)^{-1}=\begin{bmatrix}(I_{N_{0}}-A_{00})^{-1}\,,&(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\\ 0&(I_{N_{1}}-A_{\Omega\Omega})^{-1}\,.\end{bmatrix}$
In the assumed model all low-level variables are either in $\pi(0)$ or in $\Omega$. Since the CMR is constructive we have $\pi(1)\subset\Omega$. Without loss of generality, we can impose $\pi(1)=\Omega$ by setting the unused components of $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ to zero. For an arbitrary interventional setting $\mathbf{i}$, this leads to the mapping of the low-level variable to the high-level cause variable, which we denote $\widehat{Z}_{1}^{(\mathbf{i})}=\tau_{1}(\mathbf{X}^{(\mathbf{i})})$, to satisfy
$\widehat{Z}_{1}^{(\mathbf{i})}=\bm{\tau}_{1}^{\top}(I_{N}-A)^{-1}(\mathbf{U}+\mathbf{i})=\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}(\mathbf{U}_{\Omega}+\mathbf{i}_{\Omega})\,.$ (17)
Moreover, because we assume also shift interventions in the high-level model, the cause $Z_{1}$ in this model has an interventional distribution satisfying
@@
Necessary conditions. We are looking for solutions satisfying $\mathcal{L}_{\mathrm{cons}}=0$. By positivity of the KL divergence, this implies that for almost all $\mathbf{i}$, the distribution of $\widehat{Z}_{1}^{(\mathbf{i})}$ matches the learned high-level interventional distribution of high-level cause $Z_{1}$, which satisfies
Matching unintervened distributions of $Z_{1}$. If Assum. 3.2 holds, the prior $P(\mathbf{i}_{\Omega})$ has density with respect to the Lebesgue measure with support including a neighborhood of $\mathbf{i}_{\Omega}=\bm{0}$. By continuity of the KL divergence with respect to the intervention parameters, a solution making the consistency loss vanish needs to have the KL divergence term vanish for $\mathbf{i}_{\Omega}=\bm{0}$ (otherwise we could find a neighborhood of $\mathbf{i}_{\Omega}=\bm{0}$ such that the KL divergence does not vanish, by continuity of the KL divergence, and $\mathcal{L}_{\mathrm{cons}}$ would be non-vanishing, contradicting our assumption).
Alternatively, Assum. 3.3 also obviously implies vanishing of the KL divergence for the unintervened setting.
This vanishing of the KL divergence entails, again by positivity, that its terms, the two unintervened densities, are equal, such that we get, using Eq. (17)
$P^{(0)}({Z}_{1})=$ $\,(\tau_{1})_{\#}\left[P(\mathbf{X})\right]=(\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1})_{\#}\left[P(\mathbf{U}_{\Omega})\right]$ (18)
$=$ $\,\mathcal{N}(\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\bm{\mu}_{{\Omega}},\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{{\Omega}}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1})\,,$ (19)
which entails the following constraints on the variance and mean of the high-level cause
$\sigma_{Z_{1}}^{2}=\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{{\Omega}}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1}$ (20)
@@
$\mu_{Z_{1}}=\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\bm{\mu}_{{\Omega}}\,.$ (21)
Matching interventional distributions of $Z_{1}$. For the same reasons, under Assum. 3.2, we can further match the interventional distributions in an open set included in the interior of the support of $P(\mathbf{i})$, such that for all $\mathbf{i}$ in this open set the following distributions are the same
$P^{(\omega_{1}(\mathbf{i}))}({Z}_{1})=$ $\,\mathcal{N}(\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}(\bm{\mu}_{\Omega}+\mathbf{i}_{\Omega}),\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1})\,\mbox{and}$
$(\tau_{1})_{\#}\left[P^{(\mathbf{i})}(\mathbf{X})\right]=$ $\,\mathcal{N}(\mu_{Z_{1}}+{\bm{\omega}}_{1}^{\top}\mathbf{i},\sigma_{Z_{1}}^{2})=\mathcal{N}(\bar{\tau}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\bm{\mu}_{\Omega}+{\bm{\omega}}_{1}^{\top}\mathbf{i},\sigma^{2}_{Z_{1}})\,.$
Indeed, otherwise the KL would not vanish in a neighborhood of non-zero measure and would contradict the assumption that $\mathcal{L}_{\mathrm{cons}}$ vanishes.
@@
$\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}(\bm{\mu}_{\Omega}+\mathbf{i}_{\Omega})=\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\bm{\mu}_{\Omega}+\bar{{\bm{\omega}}}_{1}^{\top}\mathbf{i}_{\Omega}\,,$
@@
$\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\mathbf{i}_{\Omega}=\bar{{\bm{\omega}}}_{1}^{\top}\mathbf{i}_{\Omega}\,.$
Since this equality between two linear functions of $\mathbf{i}_{\pi(1)}$ is valid on an open set of the vector space of $\mathbf{i}_{\pi(1)}$, these functions must be equal (we can reparameterize $\mathbf{i}$ to show that the linear maps must match on a basis of the space, so they are equal). This is valid if and only if, in addition to Eqs. (20-21),
$\bar{{\bm{\omega}}}_{1}=(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1}\,,$ (22)
@@
Alternatively, we obtain the same result by replacing Assum. 3.2 by Assum. 3.3. Indeed, the finite distribution over interventions imposes that the KL term inside the expectation must vanish for each of them (including the unintervened distribution). As long as the collection of finite interventions vectors forms a rank #$\#\Omega=\#\pi(1)=N_{1}$ family, we can choose a subset of $N_{1}$ such vectors $\{\mathbf{i}^{1}_{\Omega},\dots,\mathbf{i}_{\Omega}^{N_{1}}\}$ such that it forms a linearly independent family. It can be used to build the matrix equality
$\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\left[\mathbf{i}^{1}_{\Omega},\dots,\mathbf{i}_{\Omega}^{N_{1}}\right]={\bm{\omega}}_{1}^{\top}\left[\mathbf{i}^{1}_{\Omega},\dots,\mathbf{i}_{\Omega}^{N_{1}}\right]$ (23)
where the matrix $\left[\mathbf{i}^{1}_{\Omega},\dots,\mathbf{i}_{\Omega}^{N_{1}}\right]$ is invertible. By right-multiplying Eq. (23) by this inverse, we obtain Eq. (22) again.
Matching distributions of $Y$. We can move on to check the implication of consistency of the effect’s conditional. It entails for almost all of $\mathbf{i}$
@@
Introducing $T=\begin{bmatrix}\bm{\tau}_{0}^{\top}\\ \bm{\tau}_{1}^{\top}\end{bmatrix}$ the left-hand side is obtained by using
@@
$Y=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\mathbf{U}_{\pi(0)}+\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}(\mathbf{U}_{\pi(1)}+\mathbf{i})$
@@
Given the affine mechanism assumption of Eq. (7), $f(Z)=\alpha Z+\beta$ and under Assum. 3.2, equality of marginal distributions entails the following equality for all $\mathbf{i}_{\Omega}$ in an open neighborhood of 00 (otherwise $\mathcal{L}_{\mathrm{rel}}\leq\mathcal{L}_{\mathrm{cons}}$ would not vanish)
$\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\bm{\mu}_{\pi(0)}+\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}(\bm{\mu}_{\mathbf{U}_{1}}+\mathbf{i}_{\Omega})=\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}(\mu_{\Omega}+\mathbf{i}_{\Omega})+\beta+\bm{\mu}_{R_{0}}$
@@
$\beta+\bm{\mu}_{R_{0}}=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\bm{\mu}_{{\pi(0)}}\,.$
We can fix $\bm{\mu}_{R_{0}}$ to zero to avoid redundancy of additive constants, such that
$\mu_{Y|Z=0}=\beta=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\bm{\mu}_{{\pi(0)}}$ (24)
and consistency of non-zero shift interventions additionally entail for all $\mathbf{i}_{\Omega}$ in the support of $P(\mathbf{i}_{\Omega})$
$\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\mathbf{i}_{\Omega}=\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\mathbf{i}_{\Omega}\,.$
Since one can always choose a linearly independent family of vectors $\mathbf{i}_{\Omega}$ within the open neighborhood of zero for which this equality holds, this yields
$\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}=\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\,.$
Then, right-multiplying by $(I_{N_{1}}-A_{\Omega\Omega})$, we get
$A_{0\Omega}^{\top}(I_{N_{0}}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}=\alpha\bar{\bm{\tau}}_{1}\,.$ (25)
Similarly as above, the same conclusion can be drawn if we replace Assum. 3.2 by Assum. 3.3.
Sufficiency of the constraints. We have derived the expressions of the TCR parameters in Eqs. (22,24,25) from necessary conditions for matching the marginals of high-level variables, $P(Z_{1})$ and $P(Y)$, to their corresponding pushforward distributions of the low-level variables, $(\tau_{1})_{\#}[P(\mathbf{X})]$ and $(\tau_{0})_{\#}[P(\mathbf{X})]$. Now what remains is to check the same for conditional distributions to show those conditions are sufficient. Indeed, this implies that the joint high-level distributions of $(Y,Z_{1})$ and $(\tau_{0}(\mathbf{X}),\tau_{1}(\mathbf{X}))$ are matching.
Let us first note that given that the low-level model, the $\tau$ maps and the high-level mechanisms are linear or affine, and that the low-level exogenous variables are Gaussian, the exogenous high-level variable will necessarily be Gaussian as well to satisfy the consistency constraints.
Let us now compute the covariance matrix of the low-level variables.
$cov(\mathbf{X})=\begin{bmatrix}(I_{N_{0}}-A_{00})^{-1}&(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\\ \bm{0}&(I_{N_{1}}-A_{\Omega\Omega})^{-1}\end{bmatrix}\Sigma_{U}\begin{bmatrix}(I_{N_{0}}-A_{00})^{-\top}&\bm{0}\\ (I_{N_{1}}-A_{\Omega\Omega})^{-\top}A_{0\Omega}^{\top}(I_{N_{0}}-A_{00})^{-\top}&(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\end{bmatrix}\,.$
Because the exogenous covariance is block diagonal, we get
$=\begin{bmatrix}\begin{matrix}(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}A_{0\Omega}^{\top}(I_{N_{0}}-A_{00})^{-\top}\\ +(I_{N_{0}}-A_{00})^{-1}\Sigma_{\pi(0)}(I_{N_{0}}-A_{00})^{-\top}\end{matrix}&(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\\ \hline\cr(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}A_{0\Omega}^{\top}(I_{N_{0}}-A_{00})^{-\top}&(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\end{bmatrix}\,.$
Then, if we denote $\widehat{Y}=\tau_{0}(\mathbf{X})$